OpenStudy (anonymous):

Suppose y varies directly with x, and y = 8 when x = –6. What direct variation equation relates x and y? What is the value of y when x = –2?

OpenStudy (blockcolder):

Since y varies directly as x, we have y/x=k. Since y=8 when x=-6, k=8/-6=-4/3. Thus, the equation that relates x and y is: \[\frac{y}{x}=-\frac{4}{3}\]. Now, substitute x=-2 here and you'll get its corresponding y-value.